OpenStudy (anonymous):

A girl standing on a bridge throws a stone vertically downward with an initial velocity of 15.0 m/s into the river below. If the stone hits the water 2.00 seconds later, what is the height of the bridge above the water? Dont give the answer, just tell me how to do it pls

OpenStudy (anonymous):

Well first, do you know the displacement equation in physics?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

one of newton's kinematics expressions. delta y = v(initial)x(time) + 1/2at^2; where a = gravity

OpenStudy (anonymous):

Almost - it's s = v0 t + (1/2) a t^2. the other guy forgot a t

OpenStudy (anonymous):

ninjaedit :)

OpenStudy (anonymous):

So s = v0 t + (1/2) a t^2. You know t. You know a - that's -9.81 m/s/s, the acceleration of gravity near earth's surface. You know v0 -- that's 15. Now just solve for s.

OpenStudy (anonymous):

you guys are amazing :D thank u

OpenStudy (mathmate):

Wouldn't v0=-15 (downwards!)

OpenStudy (anonymous):

ah yes, sorry. the sign of v0 should be the same as the sign for 'a'.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

s=10.38?

OpenStudy (anonymous):

i did not get that answer

OpenStudy (mathmate):

10 does not look right, 2*15m gives 30m.

OpenStudy (mathmate):

Oh, you've got a problem with the sign. Both v0 and a are negative!

OpenStudy (anonymous):

s=-15(2)+(1/2)(-9.81)(2^2) s=-30+(1/2)(-9.81)(4) s=-30+(1/2)(-39.24) s=-30+(-19.62) s=10.38

OpenStudy (mathmate):

Check the following sequence: s=-30+(-19.62) s=10.38

OpenStudy (anonymous):

the very last step is wrong. sign confusion.

OpenStudy (anonymous):

ur right. its supposed to be -49.62

OpenStudy (anonymous):

that's what i got

OpenStudy (anonymous):

but the negative doesnt make sense since its the height above water

OpenStudy (anonymous):

that's the distance from the starting point (from the bridge) to the water, with up being positive and down being negative. So the water is 49.62 feet below the bridge.

OpenStudy (anonymous):

oh ok :) sorry. my brain is fried today